1.A machine contains 5 components that all need to be functional in order for the equipment to operate. The reliability of the components is .99, .98, .99, .999, and .96, what is the resulting system reliability?

2. If the second and fourth components had redundant components with the same reliabilities what would be the system reliability?
Answer #1

1.

reliability formula :

Rsystem = R1*R2*R3*R4*R5

= 0.99*0.98*0.99*0.999*0.96

Rsystem = 0.921156

2.

if the component has redundant component with each having reliability R

Rcomponent = 1 - (1-R)^2

2nd and 4th reliability :

R2 = 1 - (1-0.98)^2 = 0.9996

R4 = 1 - (1-0.999)^2 = 0.999999

Rsystem = R1*R2*R3*R4*R5

= 0.99*0.9996*0.99*0.999999*0.96

= 0.9405187

Rsystem = 0.9405187

(please UPVOTE)